A zeolite softener 

got 85% exhausted while treating the hardness of 15,000 liters of water 

and then it requires 25 litres of 1.5gm/L NaCl solution for regeneration. 

Calculate the hardness of water sample.


1
Expert's answer
2021-02-02T01:25:13-0500

The Nacl concentration is 1.5gm/L

=> 1 ltrs of the solution contains 1.5 grms of Nacl

=> 25 ltrs of solution contains 37.5 gms of Nacl

CaCO₃ equivalent of 37.5 gms of Nacl = 37.5 x 100/(2x58.5)

= 375/117

= 3.205 gms

= 3205 mg

1 ltrs of water ------> x ppm

=> 10000 ltrs of water = x * 10000 ppm = 3205mg

=> x =3205/10000= 0.3205 ppm

Hence, the hardness of water is 0.3205 ppm